Assume that thermometer readings are normally distributed with a mean of 0 degrees °C and a standard deviation of 1.00 degrees °C. A thermometer is randomly selected and tested. For the case​ below, draw a​ sketch, and find the probability of the reading.​ (The given values are in Celsius​ degrees.) Between 0.25 and 1.75
Answer #1

SOLUTION:

From given data,

Assume that thermometer readings are normally distributed with a mean of 0 degrees °C and a standard deviation of 1.00 degrees °C. A thermometer is randomly selected and tested. For the case​ below, draw a​ sketch, and find the probability of the reading.​ (The given values are in Celsius​ degrees.) Between 0.25 and 1.75

standard deviation = 1.00

mean = 0

probability of the reading Between 0.25 and 1.75 = P(0.25 < Z <1.75 )

= P(Z < 1.75) - P(Z < 0.25)

= 0.95994 - 0.59871

= 0.36123

The sketch will be ,

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.